1.Name the type of bond that links

a.A nucleotide to another nucleotide in a DNA molecule

b. A nitrogenous base to a ribose sugar in an RNA molecule

2.Draw the structures of the amino acids tyrosine and aspartic acid


3.Briefly explain how enzymwe descrease the activation energy of biocchemical reactions


4.A nucleic acid has 13% of adenine. Determine the composition of guanine , cytosine and thymine


5.What are the forces that stabilizedouble stranded DNA?


6.Provide the structures and functions of three biochemically important disaccharides

7.What arethe structural and functional differences between starch and cellulose


1
Expert's answer
2021-05-10T17:51:02-0400

1:

(a) Phosphodiester bond. 

(b)N-Glycosidic bond. 

2:

Structure of Tyrosine:




Structure of Aspartic acid:



3: 

Enzymes decrease activation energy by attaching themselves to substrates such that bonds are changed in the substrate such that they get closer to forming transition states. This helps in decreasing the energy essential for generating a transition state. The lower the entropy of the biochemical reaction by bringing all substrate molecules together. Inside the enzyme-substrate complex they form new bonds that are not possible without enzymes; by this, they help the reaction to attain absolutely variable biochemical pathways for completion. The work by generating a charge distribution that is opposite to the transition state thus decreasing the activation energy.

4:

 Given that, 

A nucleic acid has 13% adenine. So, according to Chargaff's rule, %A =%T and % G= %C. This rule is valid for each of the two DNA strands. Hence, the % of Thymine would be 13% according to the rule.

Now, A+G = T+C 

Therefore, G+C= 100 - (2*13) =100 -26 = 74

or, G =37% and C =37%.

Hence, A=13%, T=13%, C=37%, G=37%.

5:

Forces are-  

Hydrogen bonding- present between base pairs, there are Three H-bonds between Guanine and Cytosine and Two H-bonds between Adenine and Thymine.

Van der waal forces- it is a weak force, it present between the base pairs forming a stack arrangement or we can say that it involves in base pair stacking

 Hydrophobic interaction- It provides stability and present between the nitrogenous bases and surrounding water sheath.

6:

Maltose:

It is made up of two α−D-glucose units held together by α (1 → 4)

glycosidic bond



Sucrose:

It is comprised of α−D-glucose and β−D-fructose.



  • Lactose:

It is made up of D-galactose and β-D-glucose.



7:

structural differences 

Starch comprises glucose monomers that are linked by alpha 1-4 linkages while cellulose comprises glucose monomers that are linked by beta 1-4 linkages.

The structure of starch can be linear, branched or both while the structure of cellulose is mostly linear.

 functional differences 

Starch acts as the stored form of carbohydrate in plants while cellulose serves as the main structural component of cell wall in plants.

Forms of starch are amylose and amylopectin while forms of cellulose are hemicellulose and lignin